Records taken from the number of male and 

female births in 800 families having four 

children are given below : 

No. of Births 

Frequency 

Male Female 

0 4 32 

1 3 178 

2 2 290 

3 1 236 

4 0 64 

Test whether the data is consistent with the 

hypothesis that the binomial law holds for 

the chance of a male birth is equal to that of 

a female birth at the 5% level of significance.


1
Expert's answer
2021-05-27T12:22:47-0400

Number of boys  0  1 2 3 4 


Number of girls 4 3 2 1 0


Number of families 32 178 290 236 64


P(all boys) "= (\dfrac{1}{2})^4 = \dfrac{1}{16}"


P(3 boys and 1 girl) "= ^4C_3\times (\dfrac{1}{2})^3\dfrac{1}{2} = \dfrac{1}{4}"


P(2 boys and 2 girls) "= ^4C_2\times \dfrac{1}{2}^2\times \dfrac{1}{2}^2 = \dfrac{3}{8}"


P(1 boy and 3 girls) "= ^4C_1 \times \dfrac{1}{2} \times \dfrac{1}{2}^3 = \dfrac{1}{4}"


P(all girls)"= \dfrac{1}{2}^4 = \dfrac{1}{16}"


For 800 families,


Number of Families(all boys) "= \dfrac{1}{16} \times 800 = 50"


Number of Families(3 boys and 1 girl) "= \dfrac{1}{4}\times 800 =200"


Number of Families(2 boys and 2 girls) "= \dfrac{3}{8}\times 800 = 300"


Number of Families(1 boy and 3 girls) "= \dfrac{1}{4}\times 800 = 200"


Number of Families(all girls)"= \dfrac{1}{16}\times 800 = 50"


"H_0:" The male & female births are equally probable.


Based on the information provided, the significance level is "\alpha = 0.05" the number of degrees of


freedom is "= 5-1 = 4" so then the rejection region for this test is 


"R=" { "\chi^2: \chi^2>9.488" }


The Chi-Squared statistic is computed as follows:


"x^2 = \sum \dfrac{(f_0-f_c)^2}{f_c}"


"= \dfrac{(32-50)^2}{50}+\dfrac{(178-200)^2}{200}+\dfrac{(290-300)^2}{300}+\dfrac{(236-200)^2}{200}+\dfrac{(64-50)^2}{50} = \dfrac{589}{30} = 19.633"


Since it is observed that "\chi^2 = 19.633>9.488 = \chi_c^2" it is then concluded that the null


hypothesis is rejected.


Therefore, there is enough evidence to claim that the male & female births are not equally probable, at the "\alpha = 0.05" significance level.